### v1

# Masked Image Modeling via Dynamic Token Morphing

Abstract
Masked Image Modeling (MIM) arises as a promising option for Vision Transformers among various self-supervised learning (SSL) methods. The essence of MIM lies in token-wise masked patch predictions, with targets patchified from images; or generated by pre-trained tokenizers or models. We argue targets from the pre-trained models usually exhibit spatial inconsistency, which makes it excessively challenging for the model to follow to learn more discriminative representations. To mitigate the issue, we introduce a novel self-supervision signal based on Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens. DTM can be generally applied to various SSL frameworks, yet we propose a simple MIM that employs DTM to effectively improve the performance barely introducing extra training costs. Our experiments on ImageNet-1K and ADE20K evidently demonstrate the superiority of our methods. Furthermore, the comparative evaluation of iNaturalist and Fine-grained Visual Classification datasets further validates the transferability of our method on various downstream tasks. Our code will be released publicly.
1 Introduction
Since the success of Vision Transformer (ViT), numerous training strategies have arisen for ViT [13], including self-supervised learning (SSL) methods [8, 16, 15, 6]. Recent advances in masked image modeling (MIM) [47, 17] have established it as a mainstream SSL approach for ViT. In essence, the crux of the MIM methods lies in leveraging token-wise optimization objectives derived from masked patch prediction. Consequently, MIM methods have explored various approaches to generate effective target patches, employing various pre-trained models including vision-language models [4, 28], utilizing momentum teachers [3, 51] or directly exploiting patchified images [17, 47].
However, we observe that patch representations generated by the pre-trained models tend to show spatial inconsistency (See §3). Specifically, for example, the patch-wise zero-shot classification of a pre-trained vision-language model reveals spatially inconsistent prediction results, as shown in Fig. 2. To delve deeper into the behaviors of patch representations from pre-trained models, we analyze the impacts of spatial inconsistency from the perspective of discriminability and training capability. We first observe that patch-wise ensemble prediction reveals inferior discriminability in zero-shot image classification and linear probing without token morphing, which aggregates contextually similar tokens. Furthermore, our distillation experiments reveals that token morphing without contextual awareness disrupts pre-training, whereas the model pre-trained using token morphing with context enjoys the potential to enhance representation quality. Consequently, spatially inconsistent target patches pose challenges in learning representations.
Motivated by the observations, we introduce a novel masked image modeling method with our Dynamic Token Morphing (DTM). We conjecture that training should be accelerated by the guidance of composite representations derived from morphed tokens, which aggregate contextually related tokens (see Fig. 1). More specifically, we encode the token-wise target representations and derive matching relations among tokens through DTM. The merging process is performed for both sides of tokens regarding the matching relation and aligning each morphed token with the corresponding merged target tokens. The range of our morphed tokens can vary from a single token to the whole tokens, covering from token-wise to image-level representation learning. We employ the bipartite matching algorithm for token morphing in the aspects of efficiency and efficacy.
Through extensive experiments, we verify the general applicability and scalability of our DTM. We observed that our method improves fine-tuning accuracies across various types of SSL frameworks, affirming its general applicability on SSL methods. Moreover, DTM enhances the baseline from ViT-S/16 to ViT-L/16, demonstrating its scalability. Comparison with previous SSL methods verifying its superiority. Our method shows improved transferability on the iNaturalist and Fine-grained Visual Classification datasets.
2 Related Work
Masked Image Modeling.
Inspired by the promising performance of masked language modeling, BEiT [4] successfully extends the MIM framework into the computer vision domain, using an external offline tokenizer from Dall-E [30]. iBOT [51] jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec [3] incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from multiple layers of neural networks. MAE [17] and SimMIM [47] demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, there have been several attempts to exploit the pre-trained model as a tokenizer. BEiT v2 [28] pre-trains a codebook for CLIP [29] to discretize a semantic space. MVP [43] exploits a tokenizer pre-trained with multimodal data to enhance semantics for MIM. A line of studies [44, 33] using CLIP as a teacher to use it as target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
Token Aggregation Methods.
Token aggregation has mainly been employed to enhance efficiency. Clustering is one of the commonly used token aggregation approaches. Hard clustering methods like K-Means [24], K-Medoids [1], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) enforce each data to exclusively belong to a single cluster. Bipartite matching [19] also aggregates data in a hard clustering manner, optimizing pairs of elements from two disjoint sets given objective function. Soft clustering allows data to belong to multiple clusters. LIT [27] employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods [31, 48, 35, 23] can be categorized as token aggregation methods; however, they have deeply focused on compressing tokens to achieve a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We leverage the concept of token aggregation to address spatial inconsistency in patch-level supervision, thereby enhancing the effectiveness of masked image modeling in terms of precision.
| Token morphing | Zero-shot image classification () | Linear probing () |
| 26.5 | 73.2 | |
| ✓ | 30.8 (+4.3) | 77.6 (+3.2) |
3 Motivation
In this section, we study the spatially inconsistent behavior of patch representations generated by pre-trained models and its impacts on pre-training. We visualize the spatial inconsistency of patch representations from a vision-language model in Fig. 2. Then, we conduct a quantitative analysis of the spatial inconsistency on the dataset scale in Table 1. Finally, we compare the distillation capability among token-wise, down-sampling-based, and several token aggregation-based objectives to confirm the efficacy of representation learning via context-aware token morphing.
Visualization of spatial inconsistent predictions.
To verify the spatial inconsistency among patch representations of the pre-trained model, we attempt to investigate spatially inconsistent predictions among patches through patch-wise zero-shot classification. To emphasize the spatial inconsistent representations, we conduct a comparative analysis between vanilla patch-wise prediction and prediction with smoothed tokens through token morphing, which aggregates contextually related tokens. Given the input image in Fig. 1(a), we visualize the patch-wise prediction results with and without token morphing in Fig. 1(b) and 1(c), respectively. Here, patches with accurate predictions are colored in green, while misclassified patches are colored in red. The wrong patches gradually transition to a darker red as the predicted class deviates from the ground-truth class.
| Distillation signals | Linear probing () |
| Token-wise (baseline) | 70.9 |
| Downsampling | 68.8 (-2.1) |
| Layer-wise Bipartite Matching [5] | 70.2 (-0.7) |
| Dynamic Token Morphing (Ours) | 72.2 (+1.3) |
Utilizing token morphing for predictions yields improved accuracy with 113 correct patches out of 196 and a gap of 31 patches compared to vanilla patch-wise prediction, demonstrating the importance of spatial consistency. On the other hand, despite the proximity and contextual similarity among patches, the wrong patches in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (french bulldog, gorilla, bull mastiff, hotdog, and newfoundland dog) while patches in Fig 1(c) show correct or relatively consistent predictions (french bulldog). The comparison between the two prediction results highlights the spatially inconsistent representations among patches of the pre-trained models, potentially disrupting representation learning when used as targets.
Analysis of the impact of spatial inconsistency.
To quantitatively assess the impacts of spatial inconsistency on the dataset scale, we conduct a quantitative comparison between patch-wise ensemble predictions with and without token morphing. Specifically, in the case of patch-wise ensemble prediction without token morphing, we predict the class scores for each patch and average the scores across the patches within the given image. In contrast, in the case of token morphing, we aggregate groups of semantically associated tokens and average the token representations group-wise before patch-wise ensemble prediction. We use the CLIP Base model for the encoder. We aggregate 98 tokens for token morphing, which occupy half of all tokens. Table 1 reports zero-shot image classification and linear probing results on ImageNet-1k. Patch-wise ensemble prediction with token morphing exceeds the performances without token morphing, which reveals that patch representations exhibit spatial inconsistency, and the token morphing for contextually related tokens can remedy the inconsistency.
Comparison on distillation capability varying learning objectives.
We further conduct a quantitative comparison of distillation approaches with varying configurations, as described in Fig. 3. The goal is to demonstrate the importance of resolving spatial inconsistency for representation learning while showing that indiscriminate aggregation may risk degrading the representations. We compare token-wise distillation, distillation with downsampled supervisions, distillation with layer-wise Bipartite Matching during encoding for targets, and distillation with our Dynamic Token Morphing. For each configuration, we pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K. As shown in Table 2, distillation with token morphing exceeds the performance of the token-wise distillation approach, revealing that patch representations exhibit spatial inconsistency, and token morphing for contextually related tokens can remedy this inconsistency. However, since the downsampling approach aggregates tokens without context awareness, and the layer-wise Bipartite Matching approach has a risk of harming final representations during encoding, both approaches exhibit degraded performances. The results emphasize the importance of context-aware token morphing.
4 Method
We introduce a novel masked image model training based on our Dynamic Token Morphing (DTM). In DTM, we analyze the relation between token representations and aggregate similar tokens to build morphed tokens representing various semantic information. By aligning morphed tokens from online and target encoders, the masked image model is applied in morphed token-level rather than token-level loss of conventional methods. DTM is designed for universal application across various online and target encoders. The framework of DTM is described in Fig. 4, where the DTM module is straightforwardly added to the baseline MIM scheme.
4.1 Token Encoding Process Formulation
We begin by formulating the token encoding process. Given an image , we patchify the image into N patches . We select positions of masked patches in a block-wise manner with a masking ratio so that . We mask the image patches by replacing the image patches of the position in to a shared learnable embedding . To this end, the input of the encoders becomes where for and for . The masked patches are concatenated with a learnable CLS token and fed into the online encoder with a subsequent linear head and the target encoder , and become encoded online tokens and where and , respectively.
| Framework | Method | FT. () |
| MAE + CLIP | 82.6 | |
| Feature MIM | MAE + CLIP + DTM | 83.2 (+0.6) |
| (CLIP teacher) | BEiT v2 | 84.2 |
| BEiT v2 + DTM | 84.4 (+0.2) | |
| Image-level SSL | BYOL | 81.7 |
| BYOL + DTM | 82.1 (+0.4) |
4.2 Dynamic Token Morphing
Considers -th DTM among total levels DTM, dynamic token morphing scheduling randomly assigns number of final morphed tokens and number of iterations to aggregate tokens. Then, we compute token morphing schedule , a sequence of token numbers to aggregate. Rather than samples a sequence of random numbers that satisfy , we simply define as:
At this point, the token morphing function for the -th level contextual representations can be determined by . Note that each -th level token morphing function is dynamically changed based on randomly sampled and . Eventually, we can obtain the -th level token morphing matrix , a contextual relation among tokens, from the target token representation as follows:
where indicates the aggregation of the -th token to the -th token . In addition, every token is assigned to a specific cluster, even in cases where it forms a single token cluster, and each token retains its exclusive association with a single cluster. Thus, morphing matrix should satisfies the following two conditions, and . Our DTM is universal to any token aggregation method; we simply employ bipartite matching to efficiently search for contextually similar tokens.
Based on the -th level token morphing matrix obtained from dynamic token morphing, the representations for online and target morphed tokens can be derived by:
where and denote for the representations of online and target morphed tokens, respectively.
| Method | ViT-S | ViT-B | ViT-L |
| MIM (Baseline) | 82.7 | 84.3 | 86.4 |
| DTM (Ours) | 83.1 | 85.4 | 86.7 |
| Gain | (+0.4) | (+1.1) | (+0.3) |
| Method | Pre-training epochs | Supervision | ViT-S | ViT-B | ViT-L | Seg. | |||
| ViT-S | ViT-B | ViT-L | |||||||
| Supervised models | |||||||||
| DeiT [37] | ICML 2021 | - | - | - | Label | - | 81.8 | - | - |
| DeiT-III [38] | ECCV 2022 | - | - | - | Label | - | 83.8 | 84.2 | 49.3 |
| Cosub [39] | CVPR 2023 | - | - | - | Label | 81.5 | 84.2 | 85.3 | 49.3 |
| Self-supervised models | |||||||||
| MoCo v3 [9] | ICCV 2021 | 300 | 300 | 300 | Pixel | 81.7 | 83.2 | 84.1 | 47.3 |
| BEiT [4] | ICLR 2021 | 300 | 800 | 800 | DALL-E | 81.7 | 83.2 | 85.2 | 47.1 |
| DINO [6] | ICCV 2021 | 3200 | 1600 | - | Feature | 82.0 | 83.6 | - | 46.8 |
| SplistMask [14] | arXiv 2021 | 300 | 300 | - | Pixel + Feature | 81.5 | 83.6 | - | 46.8 |
| iBOT [51] | ICLR 2022 | 3200 | 1600 | 1000 | Feature | 82.3 | 84.0 | 84.8 | 50.0 |
| MAE [17] | CVPR 2022 | - | 1600 | 1600 | Pixel | - | 83.7 | 85.6 | 48.1 |
| SimMIM [47] | CVPR 2022 | - | 800 | - | Pixel | - | 83.8 | - | - |
| MaskFeat [42] | CVPR 2022 | - | 1600 | 1600 | Feature | - | 84.0 | 85.7 | - |
| ExtreMa [45] | arXiv 2022 | - | 300 | - | Feature | - | 83.7 | - | 47.9 |
| CAE [10] | arXiv 2022 | 300 | 1600 | 1600 | Pixel + Feature | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE [18] | arXiv 2022 | - | 1600 | 1600 | Pixel + Feature | - | 84.4 | 85.3 | 50.1 |
| FD-CLIP [44] | arXiv 2022 | - | 300 | - | CLIP B/16 | - | 84.9 | - | 52.8 |
| BEiT v2 [28] | arXiv 2022 | - | 300 | 300 | CLIP B/16 | - | 85.0 | 86.6 | 52.7 |
| data2vec [3] | ICML 2022 | - | 800 | 1600 | Feature | - | 84.2 | 86.6 | - |
| mc-BEiT [22] | ECCV 2022 | - | 800 | 800 | VQGAN | - | 84.1 | 85.6 | 47.0 |
| MVP [43] | ECCV 2022 | - | 300 | 300 | CLIP B/16 | - | 84.4 | 86.3 | 52.4 |
| SdAE [11] | ECCV 2022 | - | 300 | - | Pixel | - | 84.1 | - | 48.6 |
| MSN [2] | ECCV 2022 | - | 600 | - | Feature | - | 83.4 | - | - |
| BootMAE [12] | ECCV 2022 | - | 800 | 800 | Pixel + Feature | - | 84.2 | 85.9 | 49.1 |
| CAN [25] | arXiv 2023 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
| SemMAE [21] | NIPS 2022 | - | 800 | - | Pixel | - | 83.3 | - | 46.3 |
| ConMIM [49] | ICLR 2023 | 300 | 800 | 1600 | Dictionary | 79.8 | 83.7 | 85.5 | 46.0 |
| RC-MAE [49] | ICLR 2023 | 1600 | 1600 | 1600 | Pixel | 82.0 | 83.6 | 86.1 | - |
| MixedAE [7] | CVPR 2023 | - | 1600 | - | Pixel | - | 83.9 | - | 49.8 |
| SIM [36] | CVPR 2023 | - | 1600 | - | Feature | - | 83.8 | - | - |
| HPM [41] | CVPR 2023 | - | 800 | 800 | Pixel | - | 84.2 | 85.8 | 48.5 |
| DeepMIM [33] | arXiv 2023 | - | 300 | - | CLIP B/16 | - | 84.8 | - | - |
| DTM (Ours) | - | 300 | 300 | 300 | CLIP B/16 | 83.2 | 85.4 | 86.7 | 53.1 |
4.3 Training Objective
We formulate the objective function by accumulating DTM loss, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss can be formulated as follows:
where is the i-th distance function, is a number of aggregated tokens for the -th online and target morphed tokens. Note that our DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in Eq. 4 for all the DTM losses. Therefore, the final objective function is the summation of all DTM losses, which is defined as:
5 Experiment
| Method | Birds | CUB-200 | CIFAR-10 | CIFAR-100 | Dogs | Average |
| Baseline | 87.3 (86.5±0.6) | 87.1 (86.8±0.6) | 99.2 (99.1±0.0) | 92.0 (91.9±0.3) | 86.9 (86.8±0.1) | 90.5 |
| DTM (Ours) | 88.8 (88.2±0.4) | 88.8 (88.1±0.4) | 99.3 (99.2±0.0) | 92.3 (92.1±0.2) | 87.9 (87.8±0.2) | 91.4 (+0.9) |
Pre-training setup.
We use the training dataset of ImageNet-1K for pre-training. We pre-train our framework and other SSL frameworks through DTM on ImageNet-1K. We pre-train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base model with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4. Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. We also pre-train various SSL frameworks through our DTM. Detailed pre-training setups can be found in Appendix. All models are pre-trained using 8 V100-32GB GPUs.
| Method | iNat 2018 | iNat 2019 | iNat 2021-mini |
| Baseline | 75.0 (74.6±0.6) | 81.1 (79.8±1.0) | 75.8 (75.2±0.6) |
| DTM (Ours) | 78.5 (77.4±0.6) | 81.9 (81.2±0.6) | 78.4 (77.5±0.6) |
Fine-tuning setup.
We fine-tune our pre-trained models on ImageNet-1K [34] by default. Pre-trained ViT-S, ViT-B, and ViT-L are fine-tuned for 300, 100, and 50 epochs, respectively. The pre-trained models are optimized by AdamW with a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. Additional fine-tuning details are listed in Appendix.
5.1 Effectiveness of Our Method
Applicability.
We apply DTM on BEiT v2, MAE with the CLIP teacher [29], and BYOL [15] to verify its applicability to various SSL frameworks. As shown in Table 3, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability on various frameworks beyond our baseline.
Scalability.
We further study the scalability of DTM. We compare the fine-tuning accuracies of our method to its baseline, which is pre-trained without DTM. We pre-train ViT-S/16, ViT-B/16, and ViT-L/16 with a resolution of for each method and evaluate the performance gains in fine-tuning accuracies. As shown in Table 5, DTM improved ViT-S/16, ViT-B/16, and ViT-L/16 by 0.4p, 1.1p, and 0.3p, respectively, verifying its general efficacy across different scales.
| Method | Dynamic | FT. () | |
| Baseline | - | - | 84.3 |
| DTM | ✓ | - | 84.0 (-0.3) |
| DTM | ✓ | ✓ | 85.4(+1.1) |
5.2 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K [34] datasets. The comparisons include supervised learning and SSL methods with various supervision signals. In the case when the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for fair comparison. Table 5 reports the fine-tuning accuracies of ViT-S/B/L backbones. We observe that our MIM pre-trained by DTM achieves 83.2, 85.4, 86.7 top-1 accuracy with ViT-S/16, ViT-B/16, and ViT-L/16, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP, DeepMIM, and BEiT v2 by 1.0p, 0.6 p, and 0.4p on ViT-B/16, respectively. Moreover, our method definitely outperforms other methods that leverage supervisions other than CLIP representations, which demonstrates the superiority of our method among self-supervised learning methods.
5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performances on ADE20K [50] to verify the transferability of our pre-trained model to dense prediction tasks. We follow the training and evaluation protocol [17]; the models are fine-tuned for 160K iterations using UperNet [46] with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 5 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method is effective in enhancing discriminability for dense prediction tasks.
5.4 Transfer learning
iNaturalist datasets.
We verify the improved transferability of our pre-trained model. We compare fine-tuning accuracies of the baseline and our proposed model on iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [40], which are highly imbalanced with different numbers of images per class. All the models are ViT-B/16 with a resolution of 224224. Following the protocol [20], we perform grid searches on learning rates and weight decay and report the maximum accuracy and the mean and standard deviation of the accuracies. Table 7 shows our DTM loss significantly improves the baseline with large margins, which reveals enhanced transferability.
Fine-Grained Visual Classification (FGVC) datasets.
We further validate the transferability of our method. Following the evaluation protocol as above, we conduct comparisons on FGVC datasets. Specifically, we evaluate fine-tuning accuracies on Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs through grid searches with different learning rates and weight decays. As shown in Table 6, our method outperforms the baseline overall across the datasets, which shows superior transferability and tuning robustness.
5.5 Analysis
Ablation study on the importance of dynamics.
Table 8 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ a ViT-B/16 model with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K [34]. For token morping without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 8, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, which demonstrates the importance of dynamic mechanism in DTM.
Faster convergence.
We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 4(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 4(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss compared to the baseline model, verifying better convergence. Finally, Fig. 4(c) verifies consistent effect of DTM across various pre-training epochs.
6 Conclusion
We have introduced a novel masked image modeling method with morphed tokens to address the negative impacts of spatially inconsistent target representations during pre-training. We first analyze the existence and impacts of spatial inconsistency in target representations. Specifically, we qualitatively observed spatial inconsistency among patches from pre-trained models despite proximity and contextual similarity. Then, we investigate the impacts on downstream tasks, including zero-shot image classification and linear classification. We further validate that token morphing enhances the pre-training capability of the target patches, while arbitrary aggregation, like downsampling, tends to disrupt them. The results demonstrate the necessity of token morphing to alleviate spatial inconsistency. Based on the observation, we have proposed Dynamic Token Morphing, dynamically aggregating contextually associated tokens with a random number of matching tokens. Subsequently, our method aligns the representations of morphed tokens with the representations of their corresponding morphed targets. The extensive experiments verify its general applicability, scalability, and superiority. We further validate its transferability on the iNaturalist datasets and Fine-grained Visual Classification datasets. Subsequent analysis reveals superiority over the baseline.
Limitation
Despite the potential of our method, we have verified its applicability only up to ViT-L/16. Resource limitations prevented us from performing more experiments on larger-scale models such as ViT-G. We will explore this in our future work.
Supplementary Material
This supplementary material includes additional experimental analyses of our proposed method. We first present additional examples of spatial inconsistency among patches from pre-trained models. Then, we verify the further applicability of Dynamic Token Morphing (DTM) on SSL framework beyond Masked Image Modeling methods and target models beyond CLIP [29]. Moreover, we conduct an ablation study on loss functions. Finally, we provide implementation details for both pre-training and fine-tuning on ImageNet-1K [34], as well as fine-tuning on ADE20K [50].
Appendix A More examples of spatial inconsistency
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize prediction results for each patch in the bounding boxes through patch-wise zero-shot classification with and without token morphing using CLIP [29]. Specifically, we color the patches in green or red to denote that the predicted classes are related or unrelated to the ground-truth class, respectively. Moreover, we gradually vary the lightness of the colors to exhibit the degree of similarity and dissimilarity. As shown in Fig. 6, patch representations without token morphing reveal spatially inconsistent patch-wise predictions compared to the prediction results with token morphing, which reveals the spatially inconsistent representations among patches. In addition, predictions with token morphing exhibit significantly more accurate patches compared to predictions without token morphing.
Appendix B Applicability on SLIP
We apply DTM with the SLIP [26] to verify its applicability to various target models. We compare fine-tuning accuracies of the baseline and our method pre-trained by target patches from SLIP. Specifically, we generate target tokens using SLIP along with its projector. We pre-train ViT-B/16 with a resolution of for 300 epochs and fine-tune for 100 epochs on ImageNet-1K [34] As shown in Table 9, our DTM successfully improves the fine-tuning accuracy of the baselines pre-trained with SLIP by 0.5p, which reveals its general applicability beyond CLIP [29].
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
| Case | FT. () |
| 83.2 | |
| 83.3 | |
| Smoothed | 83.2 |
| Cosine distance (Cos) | 84.2 |
| Case | Throughput (image/s) | FT. () |
| Baseline | 1460.0 | 84.3 |
| Layer-wise K-Means clustering | 1265.4 | 85.1 |
| DTM (K-Means clustering) | 489.1 | 85.4 |
| DTM (Bipartite matching) | 1446.0 | 85.4 |
Appendix C Ablation study on loss functions
We compare various loss functions for our method in Table 10. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [34] While all loss functions work well, the cosine distance works best.
| Hyperparameters | Small Size | Base Size | Large Size |
| Layers | 12 | 12 | 24 |
| Hidden size | 384 | 768 | 1024 |
| FFN inner hidden size | 1536 | 3072 | 4096 |
| Attention heads | 6 | 12 | 16 |
| Layer scale | 0.1 | 0.1 | 1e-5 |
| Patch size | |||
| Relative positional embeddings | ✓ | ||
| Shared relative positional embeddings | ✓ | ||
| Training epochs | 300 | ||
| Batch size | 1024 | ||
| Adam | (0.9, 0.98) | ||
| Base learning rate | 1.5e-4 | ||
| Learning rate schedule | Cosine | ||
| Warmup epochs | 10 | ||
| Gradient clipping | 3.0 | ||
| Dropout | ✗ | ||
| Drop path | 0 | ||
| Weight decay | 0.05 | ||
| Data Augment | RandomResizeAndCrop | ||
| Input resolution | |||
| Color jitter | 0.4 | ||
Appendix D Implementation details
Pre-training on ImageNet-1K.
The pre-training recipe for DTM mainly follows the recipe of BEiT v2 [28]. Table 12 reports the implementation details for pre-training. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16.
Fine-tuning on ImageNet-1K.
The fine-tuning recipe for DTM mainly follows the standard supervised learning recipe for ViT. Specifically, we adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively The overall recipe is detailed in Table 13.
Fine-tuning on ADE20K.
Table 14 summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K [50]. We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of . We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.
| Hyperparameters | ViT-S/16 | ViT-B/16 | ViT-L/16 |
| Fine-tuning epochs | 300 | 100 | 50 |
| Warmup epochs | 10 | 10 | 5 |
| Layer-wise learning rate decay | 0.6 | 0.6 | 0.8 |
| Batch size | 1024 | ||
| Adam | 1e-8 | ||
| Adam | (0.9, 0.999) | ||
| Base learning rate | 5e-4 | ||
| Learning rate schedule | Cosine | ||
| Repeated Aug | ✗ | ||
| Weight decay | 0.05 | ||
| Label smoothing | 0.1 | ||
| Stoch. depth | 0.1 | 0.1 | 0.2 |
| Dropout | ✗ | ||
| Gradient clipping | ✗ | ||
| Erasing prob. | 0.25 | ||
| Input resolution | |||
| Rand Augment | 9/0.5 | ||
| Mixup prob. | 0.8 | ||
| Cutmix prob. | 1.0 | ||
| Relative positional embeddings | ✓ | ||
| Shared relative positional embeddings | ✗ | ||
| Hyperparameters | ViT-B/16 |
| Input resolution | |
| Peak learning rate | 5e-5 |
| Fine-tuning steps | 160K |
| Batch size | 16 |
| Adam | 1e-8 |
| Adam | (0.9, 0.999) |
| Layer-wise learning rate decay | 0.75 |
| Minimal learning rate | 0 |
| Learning rate schedule | Linear |
| Warmup steps | 1500 |
| Dropout | ✗ |
| Stoch. depth | 0.1 |
| Weight decay | 0.05 |
| Relative positional embeddings | ✓ |
| Shared relative positional embeddings | ✗ |