### v2

# Morphing Tokens Draw Strong Masked Image Models

Abstract
Masked image modeling (MIM) is a promising option for training Vision Transformers among various self-supervised learning (SSL) methods. The essence of MIM lies in token-wise masked token predictions, with targets tokenized from images or generated by pre-trained models such as vision-language models. While tokenizers or pre-trained models are plausible MIM targets, they often offer spatially inconsistent targets even for neighboring tokens, complicating models to learn unified discriminative representations. Our pilot study confirms that addressing spatial inconsistencies has the potential to enhance representation quality. Motivated by the findings, we introduce a novel self-supervision signal called Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens to yield contextualized targets. DTM is compatible with various SSL frameworks; we showcase an improved MIM by employing DTM, barely introducing extra training costs. Our experiments on ImageNet-1K and ADE20K demonstrate the superiority of our methods compared with state-of-the-art, complex MIM methods. Furthermore, the comparative evaluation of the iNaturalists and fine-grained visual classification datasets further validates the transferability of our method on various downstream tasks. Code is available at https://github.com/naver-ai/dtm.
1 Introduction
Since the success of Vision Transformer (ViT), numerous training strategies have arisen for ViT [15], including self-supervised learning (SSL) methods [10, 20, 18, 7]. Recent advances in masked image modeling (MIM) [52, 19, 4] have established it as a mainstream SSL approach for ViT. In essence, the crux of the MIM methods lies in leveraging token-wise optimization objectives derived from predicting masked tokens to given targets. Consequently, MIM methods have explored various approaches to generate effective target tokens, employing various pre-trained models including vision-language models [5, 34], utilizing momentum teachers [4, 56] or directly exploiting patchified images [19, 52].
While tokenizers or pre-trained models have proven their potential as promising MIM targets [5, 34, 27, 49], we argue the pre-trained models tend to generate spatially noisy token representations of class information (e.g., inconsistent class labels), and their utilization as pre-training targets may incur negative impacts (See §3). For example, a pre-trained vision-language model reveals spatially inconsistent prediction results, as shown in Fig. 1. To delve deeper into the behaviors of token representations from pre-trained models, we analyze the impacts of spatially inconsistent representations from the perspective of discriminability and training capability. We initially observe reduced discriminability in zero-shot image classification and linear probing without token aggregation, also revealed with our continuous metric. Furthermore, another pilot distillation experiment reveals that the targets derived from a token aggregation method, not aware of context, disrupts pre-training. Unlikely, the model supervised by aggregated target tokens with context enjoys the potential of enhanced representation. Consequently, spatially inconsistent targets challenge learning one-to-one token maps, leading to suboptimal representation learning.
Bearing those in mind, we introduce a novel token contextualization method Dynamic Token Morphing (DTM), where token morphing is a process that links contextually similar tokens and aggregates them for coherent and comprehensive representations. We conjecture that training should be accelerated by the guidance provided by composite representations of morphed tokens, derived from the aggregation of contextually related tokens (see Fig. 3). Specifically, we encode the token-wise target representations and derive matching relations among tokens through DTM. The merging process is performed for both tokens regarding the matching relation and aligning each morphed token with the corresponding morphed target tokens. The range of morphing can vary from a single token to the entire tokens, covering from token-wise to image-level representation learning. While there are many options, we simply opt for bipartite matching for morphing, achieving both efficiency and efficacy.
Through extensive experiments, we verify our method’s general applicability and scalability. DTM improves fine-tuning accuracies as equipped with various SSL frameworks, affirming its general applicability as a basic element. Moreover, DTM consistently enhances the baselines from ViT-S/16 to ViT-L/16 with large margins, demonstrating its scalability. Comparison with previous SSL methods verifying its superiority. Our method showcases further transferability on the iNaturalist and fine-grained visual classification datasets.
2 Related Work
2.0.1 Masked image modeling.
Inspired by the promising performance of masked language modeling (MLM), BEiT [5] successfully extends MLM into the computer vision domain, using an external offline tokenizer from Dall-E [36]. iBOT [56] jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec [4] incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from the multiple layers of neural networks. MAE [19] and SimMIM [52] demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, several attempts have been made to exploit the pre-trained model as a tokenizer. BEiT v2 [34] pre-train a codebook for CLIP [35] to discretize a semantic space. MVP [48] exploits a tokenizer pre-trained with multimodal data to enhance the semantics for MIM. A line of studies [49, 38] using CLIP as a teacher to generate target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
2.0.2 Token aggregation methods.
Token aggregation can conceptually be categorized as a token clustering method and usually aims for efficiency. Hard clustering methods like K-Means [30], K-Medoids [1], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) [23] enforce each data to belong to a single cluster exclusively. Bipartite matching [24] also aggregates data in a hard clustering manner, which optimizes pairs elements from two disjoint sets given objective function. Meanwhile, soft clustering is defined to let data belong to multiple clusters. LIT [33] employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods [37, 53, 40, 28] can be categorized into token aggregation methods; however, they intensely focused on compressing tokens to aim for a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We take the concept of token aggregation to address spatially noisy target tokens in token-level supervision, thereby enhancing the efficacy of MIM in terms of precision.
3 Pilot Study
In this section, we study the behavior of token representations generated by pre-trained models and their negative impacts on pre-training.
Spatial Inconsistency.
We begin by visualizing token representations from a vision-language model [35] to illustrate spatial inconsistency.
The inconsistent class predictions observed on each patch, as depicted in Fig. 1, reveal noisy token representations.
Impact of Spatial Inconsistency.
The spatially noisy token representations has the potential to lead to an underestimation of the capabilities of the vision-language models.
To understand the extent of these effects, we quantitatively assess various metrics at a dataset scale in Table 1. The quantitative analyses reveal that enhancing spatial coherence in the token representations results in improved performance, thus highlighting the importance of addressing spatial inconsistency.
Spatial Inconsistency in Representation Learning. The lack of spatial coherence in token representations have substantial negative impacts on the representation learning. Utilizing such noisy representations as targets can undermine the capabilities of the models. We employ representation learning to assess token aggregation methods to confirm the efficacy of handling spatial inconsistency via a refined context-aware method.
| Token Aggregation | Zero-shot image classification () | Linear probing () | Averaged patch-wise cosine similarity with [CLS] |
| 26.5 | 73.2 | 0.53 | |
| ✓ | 30.8 (+4.3) | 77.6 (+3.2) | 0.56 (+0.3) |
3.1 Spatial Inconsistency
We define the spatial inconsistency among token representations with an illustrative example. Given the input image in Fig. 1(a), we visualize the token-wise predictions without and with token aggregation in Fig. 1(b) and 1(c), respectively. Correct and incorrect tokens are marked in green and red, respectively, with a gradient to darker shades of red, indicating a more significant deviation from the true class. Despite the proximity and contextual similarity among tokens, wrong tokens in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (French Bulldog, Gorilla, Bull Mastiff, Hotdog, and Newfoundland Dog) while tokens in Fig. 1(c) show correct or relatively consistent predictions (french bulldog). Moreover, token aggregation for predictions yields improved accuracy, with 113 correct tokens out of 196 and a gap of 31 tokens compared to the counterpart. The comparison between the two prediction results highlights the spatial inconsistency among tokens from a pre-trained model, which could potentially disrupt representation learning when used as targets.
3.2 Impact of Spatial Inconsistency
To quantitatively assess the impact of spatial inconsistency at a dataset scale, we compute the token-wise ensemble prediction using a global pool, with and without token aggregations. Fundamentally, we predict class scores for each token and average these scores across all tokens within the given image. When predicting with token aggregation, we group semantically relevant tokens and average their representations group-wise prior to token-wise ensemble prediction. We employ CLIP-B/16 [35] for our study, so we aggregate 98 tokens for token aggregation, which occupy half of the entire tokens.
Table 1 reports the results of zero-shot image classification, linear probing, and averaged cosine similarity between the [CLS] on ImageNet-1K. Note that we do not perform any extra training here. Our results show that using aggregated tokens consistently exceeds the performances of those without token aggregation across all metrics. This suggests that aggregating contextually related tokens can address inconsistencies and lead to performance gains.
Note that the final metric we use is an alternative: a patch-wise continuous similarity metric. We measure the cosine similarity between the [CLS] token and each patch, which is also a patch-wise metric but continuous111 We further involve a continuous metric to assess spatial inconsistency, as it tracks network responses continuously compared with discrete classification metrics.. As observed in Table 1, the averaged similarities are computed to 0.56 vs. 0.53 for each case. This new metric, directly linked to the [CLS] token, suggests a more direct relationship between improved accuracy and reduced spatial inconsistency. The trend aligns with the other discrete metric, depicting patch-wise classification of 30.8 and 26.5 with and without token aggregation, respectively. The quantitative assessments, using both continuous and discontinuous metrics, demonstrate that token representations exhibit spatial inconsistency. Taking this further, we believe more effectively addressing inconsistency would yield a greater impact.
3.3 Spatial Inconsistency in Representation Learning
We now study the impact of token aggregation on representation learning. This analysis aims to highlight the spatial inconsistency in self-distillation signal could potentially diminish the quality of representations. Our comparison ranges distillations with the signals of 1) token-wise, 2) downsampling, 3) layer-wise bipartite matching during encoding, 4) superpixel, and 5) token morphing via bipartite matching. Note that token morphing will be introduced in §4.1, but we show its effectiveness in advance. We pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K for each configuration.
First, Table 2 exhibits distillation with token aggregation methods such as token morphing via bipartite matching, and superpixel clustering outperforms that of the token-wise distillation approach. Once again, this reveals that token representations indeed exhibit spatial inconsistency, as we claimed. Furthermore, token aggregation methods for contextually related tokens could manage this inconsistency. Since the downsampling approach aggregates tokens without context awareness, and the layer-wise bipartite matching approach risks harming intermediate representations during encoding, both approaches exhibit diminished returns. The findings highlight the significance of employing context-aware token aggregations that preserve token count while addressing spatial inconsistency.
Despite superpixel clustering’s promising results, our pilot study finds that our method learning with 1) superpixel clustering and 2) context-aware bipartite matching yields accuracies of 87.1% and 87.9%, respectively, over 100 epochs trained on ImageNet-100. Therefore, we do not employ a superpixel-based clustering for our method, which is inefficient as well.
4 Method - Dynamic Token Morphing
We observed that 1) pre-trained models often produce noisy, spatial inconsistency token-level targets for learning, closely linked with performance degradation; 2) token aggregation methods could handle spatial inconsistency but are insufficient as a supervisory signal; 3) a well-designed method is needed to consider context and reduce noise more effectively. Motivated by the observations, we introduce an advanced token aggregation method Dynamic Token Morphing (DTM) for masked image modeling and self-supervised pre-training. In DTM, we contextually aggregate tokens to derive random numbers of morphed tokens aiming to encapsulate diversified semantic information. The core idea of DTM is illustrated in Fig. 3, where the DTM module is straightforwardly added to the baseline MIM scheme. DTM is demonstrated for learning conventional token-level MIMs by aligning morphed tokens from online and target encoders.
4.1 Preliminary
Token encoding.
Given an image , we patchify the image into N patches .
We select positions of masked patches in a block-wise manner with a masking ratio so that .
We mask the image patches by replacing the image patches of the position in to a learnable mask token .
To this end, the input of the encoders becomes where for and for .
The masked patches are concatenated with a learnable CLS token and fed into the online encoder with a subsequent linear head and the target encoder , and become encoded online tokens and where and , respectively.
Token morphing. Differ to token aggregation methods, which prioritize efficiency via token reduction, token morphing specifically targets the spatial inconsistency in token representations by taking context into account. It connects contextually relevant tokens to smooth them without reducing the number of tokens (i.e., maintaining the token count). We define the process with the token morphing function based on the morphing schedule , a sequence of token numbers to morph. Here, is a general notation for a function that calculates similarity using a matching algorithm (e.g., bipartite matching or K-Means clustering) and returns a token morphing matrix , where means the number of token groups after morphing. Eventually, we can morph tokens following Eq. 3.
4.2 Overall Pipeline of Dynamic Token Morphing
We present a more advanced token morphing method, Dynamic Token Morphing (DTM), here. DTM’s dynamic nature stems from its simultaneous exploration of multiple and diverse token morphing cases. This design is based on the insight that morphing numerous tokens enhances the denoising effect while morphing fewer tokens retains detailed token representations. Essentially, the diversity in the process would yield significant benefits. Since morphing a fixed number of tokens may benefit only one side, DTM balances these advantages by morphing a diverse range of tokens. Furthermore, to achieve diversified morphed tokens, we divide the morphing process into multiple processes to create multiple morphed tokens. To this end, DTM encompasses three key components: 1) dynamic scheduler for token counts, 2) token morphing via scheduler, and 3) aligning morphed tokens. The overall framework of DTM is described in Fig. 4.
4.2.1 Dynamic scheduler.
DTM generates multiple morphed tokens to ensure diversity and an extensive range of token variations. To this end, we first sample the final number of morphed tokens to remain after token morphing and the iteration number from uniform distributions, where represents the minimum number of morphed tokens and denotes the maximum number of iteration for sampling. Then, we define a token count scheduler , a sequence of token numbers that determines the number of tokens to morph for each iteration dynamically. Rather than sampling a sequence of random numbers that satisfies , we simply divide by for constant counts:
4.2.2 Token morphing via dynamic scheduler.
The token morphing function is dynamically changed based on randomly sampled and . While our DTM is universal to any token matching methods, we employ bipartite matching [24] to search for contextually similar tokens efficiently. Specifically, we split tokens into two groups, with each token in the first group matched to its closest cosine similarity counterpart in the second group.
Eventually, we obtain the token morphing matrix , a contextual relation among tokens, from the target token representations as follows:
where indicates that the token will be aggregated to the token . In addition, every token should be assigned to a specific cluster, even in cases where it forms a single token cluster, and each token should retain its exclusive association with a single cluster (i.e., and ). For the repetition , token pairs are successively matched.
Based on the token morphing matrix obtained by , the representations for online and target morphed tokens are derived by:
where and denote for the representations of online and target morphed tokens, respectively. Note that the morphed tokens are representative tokens for each token group, with their representations being smoothed specific to their respective groups.
4.2.3 Aligning morphed tokens.
We formulate the objective function by accumulating DTM losses, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss with sampled and is formulated as follows:
where is a distance function and is a number of tokens aggregated for the online and target morphed tokens. Here, we utilize to consider all tokens aggregated for the morphed tokens. Note that our DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in Eq. 4 for all the DTM losses. To further enhance the dynamism of our method, we concurrently apply multiple DTMs. Therefore, the final objective function is the summation of all DTM losses, which is defined as:
where denotes the total number of simultaneously employed DTMs.
| Method | Pre-training epochs | Supervision | ViT-S | ViT-B | ViT-L | Seg | |||
| ViT-S | ViT-B | ViT-L | |||||||
| Supervised models | |||||||||
| DeiT [42] | ICML 2021 | - | - | - | Label | - | 81.8 | - | - |
| DeiT-III [43] | ECCV 2022 | - | - | - | Label | - | 83.8 | 84.2 | 49.3 |
| Cosub [44] | CVPR 2023 | - | - | - | Label | 81.5 | 84.2 | 85.3 | 49.3 |
| Self-supervised models | |||||||||
| MoCo v3 [12] | ICCV 2021 | 300 | 300 | 300 | Pixel | 81.7 | 83.2 | 84.1 | 47.3 |
| BEiT [5] | ICLR 2021 | 300 | 800 | 800 | DALL-E | 81.7 | 83.2 | 85.2 | 47.1 |
| DINO [7] | ICCV 2021 | 3200 | 1600 | - | Feature | 82.0 | 83.6 | - | 46.8 |
| SplistMask [16] | arXiv 2021 | 300 | 300 | - | Pixel+Feat | 81.5 | 83.6 | - | 46.8 |
| iBOT [56] | ICLR 2022 | 3200 | 1600 | 1000 | Feature | 82.3 | 84.0 | 84.8 | 50.0 |
| MAE [19] | CVPR 2022 | - | 1600 | 1600 | Pixel | - | 83.7 | 85.6 | 48.1 |
| SimMIM [52] | CVPR 2022 | - | 800 | - | Pixel | - | 83.8 | - | - |
| MaskFeat [47] | CVPR 2022 | - | 1600 | 1600 | Feature | - | 84.0 | 85.7 | - |
| ExtreMa [50] | arXiv 2022 | - | 300 | - | Feature | - | 83.7 | - | 47.9 |
| CAE [11] | arXiv 2022 | 300 | 1600 | 1600 | Pixel+Feat | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE [22] | arXiv 2022 | - | 1600 | 1600 | Pixel+Feat | - | 84.4 | - | 50.1 |
| FD-CLIP [49] | arXiv 2022 | - | 300 | - | CLIP B/16 | - | 84.9 | - | 52.8 |
| BEiT v2 [34] | arXiv 2022 | - | 300 | 300 | CLIP B/16 | - | 85.0 | 86.6 | 52.7 |
| data2vec [4] | ICML 2022 | - | 800 | 1600 | Feature | - | 84.2 | 86.6 | - |
| mc-BEiT [27] | ECCV 2022 | - | 800 | 800 | VQGAN | - | 84.1 | 85.6 | 47.0 |
| MVP [48] | ECCV 2022 | - | 300 | 300 | CLIP B/16 | - | 84.4 | 86.3 | 52.4 |
| SdAE [13] | ECCV 2022 | - | 300 | - | Pixel | - | 84.1 | - | 48.6 |
| MSN [3] | ECCV 2022 | - | 600 | - | Feature | - | 83.4 | - | - |
| BootMAE [14] | ECCV 2022 | - | 800 | 800 | Pixel+Feat | - | 84.2 | 85.9 | 49.1 |
| CAN [31] | arXiv 2022 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
| SemMAE [26] | NIPS 2022 | - | 800 | - | Pixel | - | 83.3 | - | 46.3 |
| ConMIM [54] | ICLR 2023 | 300 | 800 | 1600 | Dictionary | 79.8 | 83.7 | 85.5 | 46.0 |
| RC-MAE [54] | ICLR 2023 | 1600 | 1600 | 1600 | Pixel | 82.0 | 83.6 | 86.1 | - |
| MixedAE [9] | CVPR 2023 | - | 1600 | - | Pixel | - | 83.9 | - | 49.8 |
| SIM [41] | CVPR 2023 | - | 1600 | - | Feature | - | 83.8 | - | - |
| HPM [46] | CVPR 2023 | - | 800 | 800 | Pixel | - | 84.2 | 85.8 | 48.5 |
| MIRL [21] | NeurIPS 2023 | - | 300 | 300 | Pixel | - | 84.1 | 85.4 | |
| DeepMIM [38] | arXiv 2023 | - | 300 | - | CLIP B/16 | - | 84.8 | - | - |
| dBOT [29] | ICLR 2024 | - | 1600 | 1600 | Feature | - | 84.5 | 86.6 | 49.5 |
| CrossMAE [17] | arXiv 2024 | 800 | 800 | 800 | Pixel | 79.3 | 83.7 | 85.4 | |
| DTM (ours) | - | 300 | 300 | 300 | CLIP B/16 | 83.2 | 85.4 | 86.7 | 53.1 |
5 Experiment
5.0.1 Pre-training setup.
We pre-train our framework and other SSL frameworks through DTM on ImageNet-1K. We train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base models [35] with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4 following [5, 34]. Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. For the hyperparameters of DTM, we used 2 for , 1 for , 147 for , and 14 for all . We also pre-train various SSL frameworks through our DTM with the same fundamental setups. More detailed setups for pre-training are detailed in Appendix. All models are pre-trained using 8 V100-32GB GPUs.
| Framework | Method | FT () |
| MAE + CLIP | 82.6 | |
| Feature MIM | MAE + CLIP + DTM | 83.2 (+0.6) |
| (CLIP teacher) | BEiT v2 | 84.2 |
| BEiT v2 + DTM | 84.4 (+0.2) | |
| Image-level SSL | BYOL | 81.7 |
| BYOL + DTM | 82.1 (+0.4) |
| Method | ViT-S | ViT-B | ViT-L |
| MIM (Baseline) | 82.7 | 84.3 | 86.4 |
| DTM | 83.2 | 85.4 | 86.7 |
| Gain | (+0.5) | (+1.1) | (+0.3) |
5.0.2 Fine-tuning setup.
We fine-tune our pre-trained models on ImageNet-1K [39] by default following the standard protocol [19, 34]. Pre-trained ViT-S, ViT-B, and ViT-L are fine-tuned for 300, 100, and 50 epochs, respectively. Optimization is performed with AdamW using a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. Additional fine-tuning details are listed in Appendix.
5.1 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K [39] datasets. The comparisons include supervised learning and SSL methods with various supervision signals. When the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for a fair comparison. Table 3 reports the fine-tuning accuracies of ViT-S/B/L backbones. We observe that our MIM pre-trained by DTM achieves 83.2%, 85.4%, and 86.7% top-1 accuracies with ViT-S/16, ViT-B/16, and ViT-L/16, respectively, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP [48], DeepMIM [38], and BEiT v2 [34] by 1.0p, 0.6p, and 0.4p on ViT-B/16, respectively. Moreover, our method outperforms other methods that leverage diverse supervision, demonstrating our method’s superiority among self-supervised learning methods.
5.2 Effectiveness of Our Method
5.2.1 Applicability on SSL frameworks.
We apply DTM on BEiT v2, MAE with the CLIP teacher [35], and BYOL [18] to verify its applicability to various SSL frameworks. As shown in Table 4, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability to various frameworks beyond our baseline.
5.2.2 Scalability.
We further study the scalability of DTM. We compare the fine-tuning accuracies of our method to its baseline, which is pre-trained without DTM. We pre-train ViT-S/16, ViT-B/16, and ViT-L/16 with a resolution of for each method and evaluate the performance gains in fine-tuning accuracies. As shown in Table 5, DTM improved 0.4p, 1.1p, and 0.3p for ViT-S/16, ViT-B/16, and ViT-L/16, respectively, verifying its general efficacy across the scales.
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
5.2.3 Further Applicability on SLIP.
We apply DTM with the SLIP [32] to verify its applicability to various target models. We compare fine-tuning accuracies of the baseline and our method pre-trained by target patches from SLIP. Specifically, we generate target tokens using SLIP along with its projector. We pre-train ViT-B/16 with a resolution of for 300 epochs and fine-tune for 100 epochs on ImageNet-1K [39] As shown in Table 6, our DTM successfully improves the fine-tuning accuracy of the baselines pre-trained with SLIP by 0.5p, which reveals its general applicability beyond CLIP [35].
5.2.4 Efficiency.
To verify the efficiency of our method, we compared Table 7 shows that bipartite matching is efficient and effective, significantly boosting the accuracy (+1.1%p) with only a 1% speed loss. While K-Means [30] also exhibits considerable improvements, it significantly degrades the training speed. Layer-wise K-Means shows accelerate the K-Means method by aggregating tokens within layers at the cost of degraded representations, leading to lower accuracy.
| Case | Throughput (image/s) | FT () |
| Baseline | 1458 | 84.3 |
| Layer-wise K-Means clustering | 1265 | 85.1 |
| DTM (K-Means clustering) | 489 | 85.4 |
| DTM (Bipartite matching) | 1446 | 85.4 |
5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performance on ADE20K [55] to verify the transferability of our pre-trained model. We follow the training and evaluation protocol [19]; the models are fine-tuned for 160K iterations using UperNet [51] with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 3 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method effectively enhances discriminability for dense prediction tasks.
| Method | iNat 2018 | iNat 2019 | iNat 2021-mini |
| Baseline | 75.0 (74.6±0.6) | 81.1 (79.8±1.0) | 75.8 (75.2±0.6) |
| DTM (Ours) | 78.5 (77.4±0.6) | 81.9 (81.2±0.6) | 78.4 (77.5±0.6) |
| Method | Birds | CUB-200 | CIFAR-10 | CIFAR-100 | Dogs | Average |
| Baseline | 87.3 (86.5±0.6) | 87.1 (86.8±0.6) | 99.2 (99.1±0.0) | 92.0 (91.9±0.3) | 86.9 (86.8±0.1) | 90.5 |
| DTM (Ours) | 88.8 (88.2±0.4) | 88.8 (88.1±0.4) | 99.3 (99.2±0.0) | 92.3 (92.1±0.2) | 87.9 (87.8±0.2) | 91.4 (+0.9) |
5.4 Transfer learning
5.4.1 iNaturalist datasets.
We verify the improved transferability of our pre-trained model. We compare fine-tuning accuracies of the baseline and our proposed model on iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [45], which are highly imbalanced with different numbers of images per class. All the models are ViT-B/16 with a resolution of 224224. Following the protocol [25], we perform grid searches on learning rates and weight decay and report the maximum accuracy and the mean and standard deviation of the accuracies. Table 8 shows our DTM loss significantly improves the baseline with large margins, which reveals enhanced transferability.
5.4.2 Fine-Grained Visual Classification (FGVC) datasets.
We further validate the transferability of our method. Following the evaluation protocol as above, we conduct comparisons on FGVC datasets. Specifically, we evaluate fine-tuning accuracies on Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs through grid searches with different learning rates and weight decays. As shown in Table 9, our method outperforms the baseline overall across the datasets, which shows superior transferability and tuning robustness.
| Case | Fine-tuning () |
| 84.5 | |
| 84.6 | |
| Smoothed | 84.4 |
| Cosine distance (Cos) | 84.8 |
| Case | Fine-tuning () |
| 84.7 | |
| 84.8 | |
| 84.7 |
| Method | Dynamic | FT () | |
| Baseline | - | - | 84.3 |
| DTM | ✓ | - | 84.0 (-0.3) |
| DTM | ✓ | ✓ | 85.4 (+1.1) |
5.5 Ablation study
We conduct ablation studies on loss functions, the range of token morphing steps, and target normalization. We also study the impacts of the number of morphing schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39].
5.5.1 Loss function.
We compare various options for the loss function in our method. We compared , , smoothed , and cosine distance. As shown in Table LABEL:tab:ablation_loss_ftn, the model pre-trained using cosine distance outperforms the models with other distance functions.
5.5.2 Range of token morphing iterations.
We compare fine-tuning accuracies of the pre-trained models while varying the ranges that randomly samples the number of morphing iterations in Table LABEL:tab:ablation_step. While our DTM works for all the sampling ranges, K=14 works best.
5.6 Analysis
5.6.1 Ablation study on the importance of dynamics.
Table 11 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ ViT-B/16 [15] with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K [39]. For token morphing without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 11, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, demonstrating the importance of dynamic mechanism in DTM.
5.6.2 Faster convergence of DTM.
We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 5(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 5(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss than the baseline model, verifying better convergence. Finally, Fig. 5(c) verifies the consistent effect of DTM across various pre-training epochs.
6 Conclusion
We have introduced a novel masked image modeling method based on token morphing to address the negative impacts of spatially inconsistent target representations during pre-training. We have first analyzed the existence and impacts of spatial inconsistency in target representations. Specifically, we qualitatively observed spatial inconsistency among tokens from pre-trained models despite proximity and contextual similarity. We then investigated the impacts on downstream tasks, including zero-shot/linear classification. We have further validated that context-aware token aggregation methods enhance the pre-training capability of the target tokens, while arbitrary aggregation, like downsampling, tends to disrupt them. Based on the observations, we have proposed Dynamic Token Morphing (DTM), which dynamically aggregates contextually associated tokens with a random number of matching tokens. Subsequently, our method aligns the representations of morphed tokens with the representations of their corresponding morphed targets. Our extensive experiments verified its general applicability, scalability, and superiority. We further validated its transferability, and subsequent analyses supported our method. Our study showcased that a token aggregation method could boost representation learning performance, with the potential for broader future applications and research directions.
Limitation. Despite the potential of our method, we have verified its applicability only up to ViT-L/16. Resource limitations prevented us from performing more experiments on larger-scale models such as ViT-G.
Appendix
This appendix includes additional experimental analyses of our proposed method. First, we present additional examples of spatial inconsistency among patches from pre-trained models in §0.A. Then, we validate the applicability of superpixel algorithms [2] into DTM in §0.B. We also conduct ablation studies on number of morphing schedules and effects of randomness in the number of morphing tokens, randomness in gradual token morphing, and target normalization in §0.C. Finally, §0.D provides implementation details for both pre-training and fine-tuning on ImageNet-1K [39], as well as fine-tuning on ADE20K [55].
Appendix 0.A More Examples on Spatial Inconsistency
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize token morphing results and correct maps of the patch-wise prediction results with and without token morphing using CLIP [35]. Specifically, the correct maps provide information about both correct and incorrect predictions. Patches with correct predictions are highlighted in yellow, while incorrectly predicted patches exhibit results with monotone colors. As shown in Fig. A, patch representations without token morphing reveal spatially inconsistent patch-wise predictions compared to the prediction results with token morphing, which reveals the spatially inconsistent representations among patches. In addition, predictions with token morphing exhibit significantly more correctly predicted patches than predictions without token morphing.
Appendix 0.B Compatibility with Superpixel Algorithms
We perform additional experiments to use the concept of superpixels [2, 8] in ours, both directly to tokens (layer-wise superpixel) and can be combined with DTM. The layer-wise superpixel method uses constant numbers of superpixels and iterations. Table A shows superpixel-based methods enjoy notable gains but do not exceed the bipartite matching one. While the layer-wise superpixel method also utilizes superpixels, the layer-wise token aggregation across the encoder layers bipartite matching approach risks harming intermediate representations during encoding
| Method | Fine-tuning () |
| Baseline | 79.5 |
| Layer-wise superpixel [2] | 86.7 |
| DTM (Superpixel clustering) [8] | 87.1 |
| DTM (Bipartite matching) | 87.9 |
| Method | Case | IN-1K Fine-tuning | IN-100 Fine-tuning |
| Baseline | 83.5 | 79.5 | |
| DTM | 1 | 84.8 | 87.6 |
| 2 | 84.9 | 87.9 | |
| 3 | 84.8 | 87.8 | |
| 4 | 84.8 | 87.8 |
| Method | Configurations | Fine-tuning () |
| Baseline | 79.5 | |
| DTM (ours) | Fixed number of morphing tokens | 87.4 |
| Single step morphing | 87.7 | |
| + Target normalization [19] | 87.7 | |
| Default | 87.9 |
| Hyperparameters | ViT-S/16 | ViT-B/16 | ViT-L/16 |
| Layers | 12 | 12 | 24 |
| Hidden size | 384 | 768 | 1024 |
| FFN inner hidden size | 1536 | 3072 | 4096 |
| Attention heads | 6 | 12 | 16 |
| Layer scale | 0.1 | 0.1 | 1e-5 |
| Patch size | |||
| Relative positional embeddings | ✓ | ||
| Shared relative positional embeddings | ✓ | ||
| Training epochs | 300 | ||
| Batch size | 1024 | ||
| Adam | (0.9, 0.98) | ||
| Base learning rate | 1.5e-4 | ||
| Learning rate schedule | Cosine | ||
| Warmup epochs | 10 | ||
| Gradient clipping | 3.0 | ||
| Dropout | ✗ | ||
| Drop path | 0 | ||
| Weight decay | 0.05 | ||
| Data Augment | RandomResizeAndCrop | ||
| Input resolution | |||
| Color jitter | 0.4 | ||
0.B.1 Ablation Study on the Number of Schedules
Our method can further enhance the diversity of target morphed tokens by employing multiple schedules within a single iteration. Thus, we study the effects of learning diverse morphed tokens derived from multiple morphing schedules simultaneously. We compare the fine-tuning accuracy of models pre-trained by our DTM, varying the number of schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39] and ImageNet-100 [39]. Table B reveals that exploring diverse target morphed tokens improves the representation capability of pre-trained models, leading to increased fine-tuning accuracies of at least 0.1p and 0.2p on ImageNet-1K and ImageNet-100 compared to the model pre-trained by the single DTM approach, respectively. However, the goal of experiencing diverse morphed tokens at once appears to be attained with double scheduling, resulting in no additional gains in performance through further exploration. Given that utilizing two schedules yields the best and is most efficient among all other multiple scheduling options, we adopt the double morphing scheduling approach.
Appendix 0.C Additional ablation studies
0.C.1 Ablation Study on Randomness of DTM.
0.C.1.1 Randomness in the number of morphing tokens.
Our DTM randomizes the number of morphing tokens since fixing the number of morphing tokens does not adequately generate diverse morphed tokens. To verify the effectiveness of the randomness, we compare the fine-tuning accuracies of models pre-trained using DTM with random and fixed numbers of morphing tokens. We pre-train and fine-tune the models for 100 epochs on ImageNet-100 [39]. As reported in the 2nd and 5th rows of Table C, exploring diverse morphed tokens improves the fine-tuning accuracy by 0.5p. This result demonstrates the impact of varying numbers of morphing tokens.
0.C.1.2 Randomness in gradual token morphing.
As morphed tokens can vary by the number of morphing iterations, we apply randomness in token morphing iterations. We compare fine-tuning accuracies of the pre-trained models with random and fixed iteration numbers for morphing to validate the effect of randomness. As shown in the 3rd and 5th rows of Table C, randomly selecting the iteration number for morphing enhances the performance, confirming its effectiveness.
0.C.2 Ablation Study on target normalization.
Target normalization is proven to have a significant impact on MAE [19]. Thus, we verify the impact of target normalization on our DTM. We employ ViT-B/16 with a resolution of for comparison. As shown in the 4th and 5th rows of Table C, target normalization does not yield a positive effect on our DTM, resulting in a decrease in fine-tuning accuracy from 87.9 to 87.7.
Appendix 0.D Implementation details
0.D.0.1 Pre-training on ImageNet-1K.
The pre-training recipe for DTM mainly follows the recipe of BEiT v2 [34]. Table D reports the implementation details for pre-training. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16.
0.D.0.2 Fine-tuning on ImageNet-1K.
The fine-tuning recipe for DTM mainly follows the standard supervised learning recipe for ViT. Specifically, we adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing. On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively. The overall recipe is detailed in Table E.
| Hyperparameters | ViT-S/16 | ViT-B/16 | ViT-L/16 |
| Fine-tuning epochs | 300 | 100 | 50 |
| Warmup epochs | 10 | 10 | 5 |
| Layer-wise learning rate decay | 0.6 | 0.6 | 0.8 |
| Batch size | 1024 | ||
| Adam | 1e-8 | ||
| Adam | (0.9, 0.999) | ||
| Base learning rate | 5e-4 | ||
| Learning rate schedule | Cosine | ||
| Repeated Aug | ✗ | ||
| Weight decay | 0.05 | ||
| Label smoothing | 0.1 | ||
| Stoch. depth | 0.1 | 0.1 | 0.2 |
| Dropout | ✗ | ||
| Gradient clipping | ✗ | ||
| Erasing prob. | 0.25 | ||
| Input resolution | |||
| Rand Augment | 9/0.5 | ||
| Mixup prob. | 0.8 | ||
| Cutmix prob. | 1.0 | ||
| Relative positional embeddings | ✓ | ||
| Shared relative positional embeddings | ✗ | ||
| Hyperparameters | ViT-B/16 |
| Input resolution | |
| Peak learning rate | 5e-5 |
| Fine-tuning steps | 160K |
| Batch size | 16 |
| Adam | 1e-8 |
| Adam | (0.9, 0.999) |
| Layer-wise learning rate decay | 0.75 |
| Minimal learning rate | 0 |
| Learning rate schedule | Linear |
| Warmup steps | 1500 |
| Dropout | ✗ |
| Stoch. depth | 0.1 |
| Weight decay | 0.05 |
| Relative positional embeddings | ✓ |
| Shared relative positional embeddings | ✗ |
0.D.0.3 Fine-tuning on ADE20K.
Table F summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K [55]. We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.
0.D.0.4 Transfer learning.
We follow the fine-tuning recipes for DTM to conduct transfer learning to iNaturalist datasets, including iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [45] and FGVC datasets, including Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs. However, we additionally perform grid searches of learning rates and weight decay. Specifically, we fine-tune the models with learning rates of , , and and weight decays of 0.05 and 0.1.